OpenStudy (anonymous):

evaluate \[\int\limits_{-\pi}^{\pi} f(x) dx fx = 4x^4 if -\pi \le x \le 0 or 9sinx if 0 \le x \le \pi\]

OpenStudy (anonymous):

i get -244.8157478 but is wrong or i wouldnt be askin

OpenStudy (lyrae):

Split into two integrals\[\int\limits_{-\pi}^{0} (4x ^{4}) dx + 9 \int\limits_{0}^{\pi} (\sin x) dx\]

OpenStudy (anonymous):

Don't do stuff in approximations. When you sold this integral, it will have an exact solution. If it contains pi's, e's, whatever, leave it. It is the ONLY correct answer.

OpenStudy (anonymous):

3.14 =/= pi